OpenStudy (anonymous):

For f(x)=5x=1 find F(-4)

OpenStudy (anonymous):

\[f(x)=5x+1??\]

OpenStudy (anonymous):

thats what the problem gave me

OpenStudy (anonymous):

OK ! \[f(-4)=5\times(-4)+1=-20+1=-19\]

OpenStudy (anonymous):

thank you very much i really appreciate it!!:)))